# The Geometry of Clifford Algorithms: Bernstein-Vazirani as Classical Computation in a Rotated Basis

### 12 Mar 2026 v1

@@ I Introduction: Clarifying the Origins of Parallelism
This paper proposes an alternative geometric view: the Hadamard gate is not primarily a “mixer” creating complexity, but a rotation operator reorienting the circuit components between the computational and Fourier bases (as in Chapter 3.3.4 in [7]), thereby aligning the computational frame with the oracle’s structure. In this perspective, BV belongs to a family of globally rotated Clifford circuits that remain classically simulable, as established by the Gottesman-Knill theorem (Ref. [3, Theorem 1], Ref. [4, Theorem 2]), revealing “quantum parallelism” as a basis artifact rather than an indispensable resource.
This work emphasizes hands-on step-by-step circuit transformations and simulations as first steps in building geometrical intuition behind the mathematical transformations supporting the Gottesman-Knill theorem.[3] The approach serves a two-pronged strategy: explicitly visualizing transformations while accustoming students to rigorous verification—prompting questions like ’Why do we need to verify?’, ’Where could it fail?’, and ’What aspects of our intuition may mislead us?’. These are relevant both in the sense of assuring ’Is our circuit transformation valid?’ and in the perspective of adopting a recurrential approach as mentioned in [2]. Furthermore, it introduces the concepts of “geometric rotations” and “geometric twists” in circuit design, preparing the ground for advanced discussions on how non-aligned bases between circuit components generate entanglement.
@@ III.2.2 A Note on Circuit Depth and Parallel Execution:
While the canonical BV circuit introduces $O(1)$ depth overhead due to the global $H^{\otimes n}$ basis rotations, our geometric reduction to the Z-basis reveals that this is purely an artifact of a coordinate mismatch. Because the identical $O(1)$ query complexity is achieved in the classical Z-basis formulation (which contains no superposition), we must conclude that the algorithm’s efficiency does not stem from "quantum parallelism" evaluating multiple states simultaneously. The BV circuits are always reducible to simply a sequential, classical linear evaluation—it is not a parallel algorithm in any basis. The physical depth of the computation is strictly the depth of the classical $CZ/CX$ oracle.
Nevertheless, this is allows to make two great didactic points:
@@ V Circuit Simulations in Qiskit
VI Conclusion and Outlook
By recasting the Bernstein-Vazirani algorithm as classical computation in a rotated basis, we provide a clear visual explanation of the “quantum parallelism.” The algorithm’s “exhaustive search” is shown to be a simple write operation after frame alignment.
This pedagogical reframing serves as a launchpad for more advanced concepts. The very least suggests that quantum algorithms may be constructed by starting in the computational basis and rotating them into another basis (Family II construction).
While this paper’s main part has focused on the algorithmic and structural deconstruction of the Bernstein-Vazirani circuit, the physical implications of this perspective are equally profound. Reinterpreting the oracle’s operation through the lens of continuous gauge transformations and fractional translations offers a deeper physical intuition into why such circuits are classically simulable. To put emphasis on the consequences of this approach, in Appendix A we provide an example in deriving the irreducible form of a Deutsch-Jozsa circuit (itself being on the boundary of Family II / III circuits).
@@ VI Conclusion and Outlook
VI.1 The Standard View: Complex Entanglement
Consider a 4-qubit DJ oracle designed to mark exactly four states (e.g., $1\rangle, |100|1$$0\rangle$, $|101\rangle$, and $|110\rangle$). In the standard computational basis, this oracle is constructed using a sequence of Multi-Controlled-X (Toffoli) gates wrapped in $X$ gates to trigger on specific bit-patterns.
@@ VI.1 The Standard View: Complex Entanglement
VI.2 Global Rotation and Target Severing
@@
VI.3 The Algebraic Collapse of the Phase Polynomial
@@
1. Gate push-through: For this particular case, we choose the $X$ on the q3 line and push it through the $CCZ$ gate in the second string block, resulting in an $XX=I$ removal and a new $CZ$ gate - see Fig. 9(a). We follow by pushing the resultant $CZ$ gate to the leftmost side of the oracle, past the $X$ gates and past the $CCZ$ using its commutativity. This results in spawning two more $Z$ gates - see Fig. 9(b). These $Z$ gates reduce via $ZZ=I$ leaving only $CCZ$ gates with accompanying $X$ gates. An interesting topic to show students is the “generation” of twists (spawning gates) due to these transformations!
@@
VI.4 The Irreducible Form and Future Work
@@
$U_{\text{oracle}}=CZ_{0,1}\cdot Z_{2}$ (4)
@@
$U_{\text{oracle}}=RXX_{0,1}\cdot RX\otimes RX\otimes X_{2}$ (5)
@@
What initially appeared to be a complex, highly entangled quantum search is revealed in the rotated basis to be a trivial evaluation of the phase polynomial $f(q)=(q_{0}\cdot q_{1})\oplus q_{2}$. The somewhat “mysterious” quantum parallelism is more simply explained as classical linear algebra evaluated in the conjugate Fourier coordinate system, with the genuine quantum entanglement originating entirely from the non-aligned topological twist of the Mølmer–Sørensen $RXX$ gate. This is a further example of possible benefits of teaching about parity gates and the Cartan decomposition (see [14, 15]). Furthermore, by demonstrating the equivalence of circuits in both the X and Z bases, the students’ attention may be redirected toward the truly indispensable quantum resource: entanglement.

### 17 Mar 2026 v2

@@ I Introduction: Clarifying the Origins of Parallelism
This paper proposes an alternative geometric view: the Hadamard gate is not primarily a “mixer” creating complexity, but a rotation operator reorienting the circuit components between the computational and Fourier bases (as in Chapter 3.3.4 in [7]), thereby aligning the computational frame with the oracle’s structure. In this perspective, BV belongs to a family of globally rotated Clifford circuits that remain classically simulable, as established by the Gottesman-Knill theorem (Ref. [3, Theorem 1], Ref. [4, Theorem 2]), revealing “quantum parallelism” as a basis-dependent pheonomenon rather than an indispensable resource.
This work emphasizes hands-on step-by-step circuit transformations and simulations as first steps in building geometrical intuition behind the mathematical transformations supporting the Gottesman-Knill theorem [3]. The approach serves a two-pronged strategy: explicitly visualizing transformations while accustoming students to rigorous verification—prompting questions like ’Why do we need to verify?’, ’Where could it fail?’, and ’What aspects of our intuition may mislead us?’. These are relevant both in the sense of assuring ’Is our circuit transformation valid?’ and in the perspective of adopting a recurrential approach as mentioned in [2]. Furthermore, it introduces the concepts of “geometric rotations” and “geometric twists” in circuit design, preparing the ground for advanced discussions on how non-aligned bases between circuit components generate entanglement.
@@ III.2.2 A Note on Circuit Depth and Parallel Execution:
While the canonical BV circuit introduces $O(1)$ depth overhead due to the global $H^{\otimes n}$ basis rotations, our geometric reduction to the Z-basis reveals that this is purely an artifact of a coordinate mismatch. While the standard algorithmic definition of quantum parallelism—evaluating a function over a coherent superposition of computational-basis states—remains valid for the standard BV circuit, our reduction demonstrates that this parallelism can be entirely bypassed via geometric transformations. Because the identical $O(1)$ query complexity is naturally recovered in the conjugate Fourier basis (which contains no superposition), educators can offer a concrete, parallel-free alternative: one can either evaluate a superposition of inputs using phase kickback and interference in the Z-basis, or treat the initial and final Hadamards as transforming the oracle itself into a classical operation in the X-basis. Ultimately, whether one observes simultaneous evaluation or a classical sequential query depends entirely on the chosen coordinate lens. The physical depth of the computation remains strictly the depth of the classical $CZ/CX$ oracle.
Nevertheless, this allows us to make two great didactic points:
@@ V Circuit Simulations in Qiskit
VI Limitations and Future Work
While the geometric reframing presented in this paper offers a rigorous circuit equivalence, it is primarily proposed as a theoretical pedagogical framework. Its practical efficacy in preventing student misconceptions compared to the standard interference-based narrative remains to be verified through formal Physics Education Research (PER) in a classroom setting. Furthermore, this approach does not attempt to resolve fundamental ontological debates regarding the physical reality of superposition; it merely utilizes basis transformations to optimize the representation of the circuit for teaching purposes.
Future work will naturally extend this geometric taxonomy beyond the globally rotated Family II circuits. A forthcoming paper will explore how algorithms introducing non-linear boolean terms—such as the quadratic Deutsch-Jozsa oracle—geometrically fracture the Pauli normalizer. Analyzing these topologies will visually demonstrate the transition from classically simulable Clifford entanglement to universal quantum computation.
@@ VI Limitations and Future Work
VII Conclusion
By recasting the Bernstein-Vazirani algorithm as a classical computation in a rotated basis, we provide a visual framework that allows educators and students to view “quantum parallelism” as a basis-dependent phenomenon. Seen through the lens of the Fourier basis, the algorithm’s apparent “exhaustive search” over a superposition transforms into a deterministic write operation after proper frame alignment.
This pedagogical reframing serves as a launchpad for more advanced concepts. At the very least, it suggests that quantum algorithms may be constructed by starting with a classical circuit in the computational basis and rotating it into another basis (a Family II construction).
While this paper’s main focus has been on the algorithmic and structural deconstruction of the Bernstein-Vazirani circuit, the physical implications of this perspective are equally profound. Reinterpreting the oracle’s operation through the lens of continuous gauge transformations and fractional translations offers a deeper physical intuition into why such circuits are classically simulable. To emphasize the consequences of this approach, in Appendix A we provide an example of deriving the irreducible form of a Deutsch-Jozsa circuit (which itself sits on the boundary between Family II and Family III circuits).
@@ VII Conclusion
VII.1 The Standard View: Complex Entanglement
Consider a 4-qubit DJ oracle designed to mark exactly four states (e.g., $1\rangle, |100|1$$0\rangle$, $|101\rangle$, and $|110\rangle$ - in Qiskit notation). In the standard computational basis, this oracle is constructed using a sequence of Multi-Controlled-X (Toffoli) gates wrapped in $X$ gates to trigger on specific bit-patterns.
@@ VII.1 The Standard View: Complex Entanglement
VII.2 Global Rotation and Target Severing
@@
VII.3 The Algebraic Collapse of the Phase Polynomial
@@
1. Gate push-through: For this particular case, we choose the $X$ on the $q_{2}$ line and push it through the $CCZ$ gate in the second string block, resulting in an $XX=I$ removal and a new $CZ$ gate - see Fig. 9(a). We follow by pushing the resultant $CZ$ gate to the leftmost side of the oracle, past the $X$ gates and past the $CCZ$ using its commutativity. This results in spawning two more $Z$ gates - see Fig. 9(b). These $Z$ gates reduce via $ZZ=I$ leaving only $CCZ$ gates with accompanying $X$ gates. An interesting topic to show students is the “generation” of twists (spawning gates) due to these transformations!
@@
VII.4 The Irreducible Form and Future Work
@@
$U_{\text{oracle}}=CZ_{0,1}\otimes Z_{2}$ (4)
@@
$U_{\text{oracle}}=RXX_{0,1}\cdot RX_{0}\otimes RX_{1}\otimes X_{2}$ (5)
@@
What initially appeared to be a complex, highly entangled quantum search is revealed in the rotated basis to be a trivial evaluation of the phase polynomial $f(q)=(q_{0}\cdot q_{1})\oplus q_{2}$. The somewhat “mysterious” quantum parallelism is more simply explained as classical algebra evaluated in the conjugate Fourier coordinate system, with the genuine quantum entanglement originating entirely from the non-aligned topological twist of the Mølmer–Sørensen $RXX$ gate. While the Bernstein-Vazirani oracle (Family II) operates entirely as linear boolean algebra, the Deutsch-Jozsa oracle may introduce non-linearity (e.g., the quadratic term above). This is a further example of possible benefits of teaching about parity gates and the Cartan decomposition (see [14, 15]). Furthermore, by demonstrating the equivalence of circuits in both the X and Z bases, the students’ attention may be redirected toward the truly indispensable quantum resource: entanglement.